# Gravitational Wave Spectral Shapes as a probe of Long Lived Right-handed Neutrinos, Leptogenesis and Dark Matter: Global versus Local BLB-L Cosmic Strings

### v1

@@ Abstract
The scale of the seesaw mechanism is typically much larger than the electroweak scale. This hierarchy can be naturally explained by $U(1)_{B-L}$ symmetry, which after spontaneous symmetry breaking, simultaneously generates Majorana masses for neutrinos and produces a network of cosmic strings. Such strings generate a gravitational wave (GW) spectrum which is expected to be almost uniform in frequency unless there is a departure from the usual early radiation domination. We explore this possibility in Type I, II and III seesaw frameworks, finding that only for Type-I, long-lived right-handed neutrinos (RHN) may provide a period of early matter domination for parts of the parameter space, even if they are thermally produced. Such a period leaves distinctive imprints in the GW spectrum in the form of characteristic breaks and a knee feature, arising due to the end and start of the periods of RHN domination. These features, if detected, directly determine the mass $M$, and effective neutrino mass $\tilde{m}$ of the dominating RHN. We find that GW detectors like LISA and ET could probe RHN masses in the range $M\in[0.1,10^{9}]$ GeV and effective neutrino masses in the $\tilde{m}\in[10^{-10},10^{-8}]$ eV range. We investigate the phenomenological implications of long-lived right-handed neutrinos for both local and global $U(1)_{B-L}$ strings, focusing on dark matter production and leptogenesis. We map the viable and detectable parameter space for successful baryogenesis and asymmetric dark matter production from right-handed neutrino decays. We derive analytical and semi-analytical relations correlating the characteristic gravitational-wave frequencies to the neutrino parameters $\tilde{m}$ and $M$, as well as to the relic abundances of dark matter and baryons. We find that the detectable parameter space reaches the boundary of hierarchical leptogenesis and encompasses a substantial portion of the near-resonant regime.
@@ I Introduction
Phase transitions Mazumdar_2019 ; Athron:2023xlk ; quiros1999finitetemperaturefieldtheory associated with the spontaneous breaking of global or gauged $U(1)_{B-L}$ symmetries generically lead to the formation of cosmic strings Vilenkin:2000jqa . These one-dimensional topological defects persist after formation and emit gravitational radiation through loop production and decay, sourcing a stochastic gravitational wave background (GWB) Fu:2023nrn ; King:2020hyd ; Ghoshal:2023sfa ; Dror:2019syi . The spectrum of this background is sensitive not only to the symmetry-breaking scale but also to the intervening cosmological history, making string-sourced GWBs a powerful probe of early Universe dynamics Cui:2017ufi ; Cui:2018rwi ; Gouttenoire:2019kij ; Gouttenoire:2019rtn ; Blasi:2020mfx ; Ghoshal:2023sfa ; Datta:2025yow .
@@
In standard scenarios, the string network evolves during radiation domination Cui:2018rwi and the resulting gravitational wave spectrum is almost uniform over many decades of frequency. Any departure from such a flat spectrum speaks to a surprise in our cosmic history Cui:2018rwi ; Ghoshal:2023sfa ; Blasi:2020wpy ; Ferrer:2023uwz with a period of early matter domination providing a particularly striking feature in the otherwise flat spectrum Cui:2017ufi ; Ghoshal:2023sfa ; Datta:2020bht ; Samanta:2021zzk ; Chianese:2024gee ; Datta:2024bqp . Specifically, the transient era of matter domination imprints two distinct high-frequency features: a transition from a flat to a power-law spectrum and a knee arising from the superposition of modes during the transition. Together, they encode the onset and duration of the matter-dominated phase, revealing information about its physical origin. With so many decades of frequency probed by current and planned gravitational wave detectors Aasi:2014mqd ; Yagi:2011wg ; Punturo:2010zz ; Hild:2010id ; Evans:2016mbw ; Sesana:2019vho , there is a lot of opportunity for one or both observables to be detected in the foreseeable future. A period of early matter domination can be caused by a metastable, long-lived particle Coughlan:1983ci ; Starobinsky:1994bd ; Dine:1995uk ; Moroi:1999zb ; Ghoshal:2022ruy . While scalar fields are often assumed to be responsible for such matter domination, in $U(1)_{B-L}$ extended seesaw models, an alternative arises naturally: the heavy right-handed neutrinos responsible for neutrino mass generation can dominate the energy density before decaying Berbig:2023yyy . In this paper, we investigate such scenarios in detail and show this can happen for the extended type I seesaw Minkowski:1977sc ; Gell-Mann:1979vob ; Yanagida:1979as ; Mohapatra:1980yp , but not type II Magg:1980ut ; Wetterich:1981bx ; Schechter:1980gr or type III Foot:1988aq ; Ma_1998 ; Ma_2002 . We further derive numerical relations linking these GW spectral features to the mass and effective neutrino mass of the right-handed neutrino responsible for the transient matter-dominated era.
@@
The decays of these right-handed neutrinos are closely tied to the origin of the baryon asymmetry via leptogenesis Fukugita:1986hr . In Leptogenesis out-of-equilibrium, CP-violating decays of right-handed neutrinos can generate a lepton asymmetry which is subsequently converted into a baryon asymmetry via electroweak sphalerons Luty:1992un ; Giudice_2004 ; Covi_1996 ; Buchm_ller_2005 ; Sakharov:1967dj ; Kolb:1979qa ; Khlebnikov:1988sr . In the presence of an intermediate matter-dominated phase, this mechanism is altered in two essential ways. First, the expansion rate is modified by matter domination, changing the dynamics of lepton asymmetry generation. Second, the entropy injected by right-handed neutrino decays dilutes the resulting asymmetry, making the final baryon-to-entropy ratio sensitive to the duration and timing of decay. Finally, right-handed neutrino decays can also furnish a dark matter production mechanism Falkowski_2011 ; Barman_2022 , which can result in either symmetric or asymmetric dark matter, and we explore both possibilities. A schematic overview of the framework is presented in Figure 1.
@@
Figure 1: The $U(1)_{B-L}$ symmetry breaking generates heavy seesaw states and an initial gravitational-wave background. These heavy states can induce a period of matter domination; in particular, only the Type-I seesaw, where the heavy seesaw state (the right-handed neutrino) dominates the energy density, can achieve such matter domination. This leads to a modified gravitational-wave background (GWB) spectral shape, while the subsequent decays of the right-handed neutrinos can explain the baryon asymmetry via leptogenesis or, alternatively and simultaneously, produce dark matter.
This paper is organised as follows: In section II, we derive the conditions under which right-handed neutrinos dominate the energy density in both global and gauged BLB-L scenarios, and show that such domination occurs only in the Type I seesaw. We then derive relations for the onset and duration of matter domination. Section III analyses the impact of this matter-dominated epoch on the gravitational wave spectrum from cosmic strings. Section IV examines the consequences for baryogenesis, providing an analytic and numerical treatment of lepton asymmetry generation and the lower bound on right-handed neutrino mass for successful leptogenesis. Section V explores the implications for dark matter production from right-handed neutrino decays. We conclude in Section VI.
@@ II Early Matter Domination in $U(1)_{B-L}$ Seesaw
The seesaw mechanism naturally accounts for the smallness of active neutrino masses by introducing heavy states: fermion singlets (right-handed neutrinos) in type-I, scalar triplets in type-II, or fermion triplets in type-III. The masses of these heavy states are inversely proportional to the active neutrino masses Minkowski:1977sc ; Gell-Mann:1979vob ; Yanagida:1979as ; Mohapatra:1980yp ; Magg:1980ut ; Wetterich:1981bx ; Schechter:1980gr ; Foot:1988aq ; Ma_1998 ; Ma_2002 ,
@@
where $M$ is the mass of the heavy seesaw state and $v_{H}=246$ GeV is the Higgs vacuum expectation value (vev). A large $M$ naturally explains the smallness of active neutrino masses. A $U(1)_{B-L}$ symmetry provides a natural explanation for why $M$ is so large: the heavy seesaw masses are tied to the scale at which BLB-L is spontaneously broken, $v_{B-L}$, so if the scale of symmetry breaking is large, then so is the mass of the heavy seesaw state. Moreover, $U(1)_{B-L}$ is not always ad hoc, but appears in many Grand Unified Theory symmetry-breaking chains, giving further motivation for embedding the seesaw mechanism within this framework Gell-Mann:1979vob ; Buchmuller:2005eh ; Dror:2019syi ; Dunsky:2021tih . The particle content for each type of seesaw is shown in Table 1.
@@ II.1 The Condition for Early Matter Domination
In this section, we derive the analytic conditions for right-handed neutrino matter domination. We first present an approximate derivation, following the standard approach in the literature. While convenient, this estimate is inaccurate by roughly half an order of magnitude. We therefore provide a more accurate derivation, which can still be carried out analytically.
@@ II.1.1 Derivation I: Canonical derivation
As a concrete example, the Type-I seesaw introduces Standard Model singlets $N_{i}$ with degrees of freedom $g=\tfrac{7}{8}\times 2$. This gives Giudice_1999
@@
This illustrates how the onset of matter domination in seesaw scenarios is directly controlled by the heavy state mass.
@@ II.1.2 Derivation II: Comoving Abundance Method
Once $T\ll M$, the species behaves as non-relativistic matter and its energy density is well approximated by $\rho_{N}=M\,Y\,s(T)$. In this regime, the average particle energy, averaging over momentum space, is $\langle E\rangle\simeq M+\mathcal{O}(T)$, so thermal corrections are negligible compared to the mass. Equating the right-handed neutrino and radiation energy densities $\rho_{N}=\rho_{R}$, sets the temperature for matter domination
@@
where the superscript $i$ denotes the initial value. In what follows, we adopt Derivation II. For successful matter domination by such particles, the following three conditions must be satisfied:
@@
    The particle must not have decayed by the time of matter domination:
@@
We now investigate these three conditions for the $U(1)_{B-L}$ extended seesaw frameworks. If instead we consider non-thermal production channels, such as inflaton decay Asaka_1999 ; Hahn_Woernle_2009 ; Ghoshal:2022fud , curvaton decay Fong:2023egk , generic and modulated sneutrino decay Mazumdar:2003va ; Allahverdi:2002gz ; Mazumdar:2003bs ; Afzal:2024hwj , Q-ball decay Fujii:2001sn , from phase transition bubbles Dasgupta:2022isg ; Cataldi:2024pgt , preheating Cui:2024vws and reheating Ghoshal:2022kqp , or primordial black holes evaporations Datta:2020bht ; das2021lowscaleleptogenesisdark ; Perez_Gonzalez_2021 ; Barman_2022 ; Bernal_2022 ; Ghoshal:2023sfa , only conditions (ii) and (iii) need to be imposed. To remain general, we shall leave the initial abundance $Y_{i}$ as a free parameter throughout our analysis.
@@ II.2.1 Type I: Global $U(1)_{B-L}$ case
The type I seesaw mechanism Minkowski:1977sc ; Gell-Mann:1979vob ; Yanagida:1979as is among the most economical and widely studied explanations for the origin of light neutrino masses. In this framework, heavy right-handed neutrinos $N_{i}$ are introduced, which couple to the Standard Model lepton doublets $L_{\alpha}$ and the Higgs doublet $H$ through Yukawa interactions. After electroweak symmetry breaking, these interactions generate a Dirac mass term for the neutrinos. If the right-handed neutrinos also acquire large Majorana masses, the light neutrinos obtain naturally small masses via the seesaw mechanism. Since the Majorana mass term explicitly violates BLB-L by two units, the global $U(1)_{B-L}$ symmetry requires the introduction of an additional complex scalar field $\Phi$, carrying non-zero BLB-L charge Fu:2023nrn . When $\Phi$ acquires a vacuum expectation value, the BLB-L symmetry is broken, the right-handed neutrinos obtain Majorana masses, and the seesaw mechanism is realized. The Lagrangian is augmented by Yukawa, Majoron and Higgs portal interactions, as well as a quartic potential for the symmetry-breaking scalar
@@
$\Gamma_{\text{prod}}\sim\left(\frac{\lambda_{\phi}\ v_{B-L}\ \tilde{y}_{i}}{m_{\phi}^{2}}\right)^{2}T,\quad\Rightarrow\quad\frac{\Gamma}{H}\sim\left(\frac{\lambda_{\phi}\ v_{B-L}\ \tilde{y}_{i}}{m_{\phi}^{2}}\right)^{2}\ \frac{M_{\text{Pl}}}{T}\ .$ (13)
@@
$\frac{\Gamma}{H}\propto\frac{1}{T^{5}}$ (14)

### v2

@@ Abstract
The scale of the seesaw mechanism is typically much larger than the electroweak scale. This hierarchy can be naturally explained by $U(1)_{B-L}$ symmetry, which after spontaneous symmetry breaking, simultaneously generates Majorana masses for neutrinos and produces a network of cosmic strings. Such strings generate a gravitational wave (GW) spectrum which is expected to be almost uniform in frequency unless there is a departure from the usual early radiation domination. We explore this possibility in Type I, II and III seesaw frameworks, finding that only for Type-I, long-lived right-handed neutrinos (RHN) may provide a period of early matter domination for parts of the parameter space, even if they are thermally produced. Such a period leaves distinctive imprints in the GW spectrum in the form of characteristic breaks and a knee feature, arising due to the end and start of the periods of RHN domination. These features, if detected, directly determine the right-handed neutrino mass $M$, and associated left-handed effective neutrino mass $\tilde{m}$ of the dominating RHN. We find that GW detectors like LISA and ET could probe RHN masses in the range $M\in[0.1,10^{9}]$ GeV and effective neutrino masses in the $\tilde{m}\in[10^{-10},10^{-8}]$ eV range. We investigate the phenomenological implications of long-lived right-handed neutrinos for both local and global $U(1)_{B-L}$ strings, focusing on dark matter production and leptogenesis. We map the viable and detectable parameter space for successful baryogenesis and asymmetric dark matter production from right-handed neutrino decays. We derive analytical and semi-analytical relations correlating the characteristic gravitational-wave frequencies to the neutrino parameters $\tilde{m}$ and $M$, as well as to the relic abundances of dark matter and baryons. We find that the detectable parameter space reaches the boundary of hierarchical leptogenesis and encompasses a substantial portion of the near-resonant regime.
@@ I Introduction
Phase transitions Mazumdar and White (2019); Athron et al. (2024); Quiros (1999) associated with the spontaneous breaking of global or gauged $U(1)_{B-L}$ symmetries generically lead to the formation of cosmic strings Vilenkin and Shellard (2000). These one-dimensional topological defects persist after formation and emit gravitational radiation through loop production and decay, sourcing a stochastic gravitational wave background (GWB) Fu et al. (2023); King et al. (2021); Ghoshal et al. (2023a); Dror et al. (2020). The spectrum of this background is sensitive not only to the symmetry-breaking scale but also to the intervening cosmological history, making string-sourced GWBs a powerful probe of early Universe dynamics Cui et al. (2018, 2019); Gouttenoire et al. (2020a, b); Blasi et al. (2021); Ghoshal et al. (2023a); Datta et al. (2025); Roshan and White (2025).
@@
In standard scenarios, the string network evolves during radiation domination Cui et al. (2019) and the resulting gravitational wave spectrum is almost uniform over many decades of frequency. Any departure from such a flat spectrum speaks to a surprise in our cosmic history Cui et al. (2019); Ghoshal et al. (2023a); Blasi et al. (2020); Ferrer et al. (2023) with a period of early matter domination providing a particularly striking feature in the otherwise flat spectrum Cui et al. (2018); Ghoshal et al. (2023a); Datta et al. (2021); Samanta and Datta (2021); Chianese et al. (2025); Datta and Samanta (2025a). Specifically, the transient era of early matter domination imprints two distinct high-frequency features: a transition from a flat to a power-law spectrum and a knee arising from the superposition of modes during the transition. Together, they encode the onset and duration of the matter-dominated phase, revealing information about its physical origin. With so many decades of frequency probed by current and planned gravitational wave detectors Aasi and others (2015b); Yagi and Seto (2011); Punturo and others (2010); Hild and others (2011); Abbott and others (2017d); Sesana and others (2021), there is a lot of opportunity for one or both observables to be detected in the foreseeable future. A period of early matter domination can be caused by a metastable, long-lived particle Coughlan et al. (1983); Starobinsky and Yokoyama (1994); Dine et al. (1995); Moroi and Randall (2000); Ghoshal et al. (2022). While scalar fields are often assumed to be responsible for such early matter domination, in $U(1)_{B-L}$ extended seesaw models, an alternative arises naturally: the heavy right-handed neutrinos responsible for neutrino mass generation can dominate the energy density before decaying Berbig and Ghoshal (2023). In this paper, we investigate such scenarios in detail and show this can happen for the extended type I seesaw Minkowski (1977); Gell-Mann et al. (1979); Yanagida (1979); Mohapatra and Senjanovic (1981), but not type II Magg and Wetterich (1980); Wetterich (1981); Schechter and Valle (1980) or type III Foot et al. (1989); Ma (1998); Ma and Roy (2002). We further derive numerical relations linking these GW spectral features to the mass and effective neutrino mass of the right-handed neutrino responsible for the transient matter-dominated era.
@@
The decays of these right-handed neutrinos are closely tied to the origin of the baryon asymmetry via leptogenesis Fukugita and Yanagida (1986). In Leptogenesis out-of-equilibrium, CP-violating decays of right-handed neutrinos can generate a lepton asymmetry which is subsequently converted into a baryon asymmetry via electroweak sphalerons Luty (1992); Giudice et al. (2004); Covi et al. (1996); Buchmüller et al. (2005); Sakharov (1967); Kolb and Wolfram (1980); Khlebnikov and Shaposhnikov (1988). In the presence of an intermediate matter-dominated phase, this mechanism is altered in two essential ways. First, the expansion rate is modified by early matter domination, changing the dynamics of lepton asymmetry generation. Second, the entropy injected by right-handed neutrino decays dilutes the resulting asymmetry, making the final baryon-to-entropy ratio sensitive to the duration and timing of decay. Finally, right-handed neutrino decays can also furnish a dark matter production mechanism Falkowski et al. (2011); Barman et al. (2022), which can result in either symmetric or asymmetric dark matter, and we explore both possibilities. A schematic overview of the framework is presented in Figure 1.
@@
Figure 1: The $U(1)_{B-L}$ symmetry breaking generates heavy seesaw states and an initial gravitational-wave background. These heavy states can induce a period of early matter domination; in particular, only the Type-I seesaw, where the heavy seesaw state (the right-handed neutrino) dominates the energy density, can achieve such early matter domination. This leads to a modified gravitational-wave background (GWB) spectral shape, while the subsequent decays of the right-handed neutrinos can explain the baryon asymmetry via leptogenesis or, alternatively and simultaneously, produce dark matter.
This paper is organised as follows: In section II, we derive the conditions under which right-handed neutrinos dominate the energy density in both global and gauged BLB-L scenarios, and show that such domination occurs only in the Type I seesaw. We then derive relations for the onset and duration of early matter domination. Section III analyses the impact of this matter-dominated epoch on the gravitational wave spectrum from cosmic strings. Section IV examines the consequences for baryogenesis, providing an analytic and numerical treatment of lepton asymmetry generation and the lower bound on right-handed neutrino mass for successful leptogenesis. Section V explores the implications for dark matter production from right-handed neutrino decays. We conclude in Section VI.
@@ II Early Matter Domination in $U(1)_{B-L}$ Seesaw
The seesaw mechanism naturally accounts for the smallness of active neutrino masses by introducing heavy states: fermion singlets (right-handed neutrinos) in type-I, scalar triplets in type-II, or fermion triplets in type-III. The masses of these heavy states are inversely proportional to the active neutrino masses Minkowski (1977); Gell-Mann et al. (1979); Yanagida (1979); Mohapatra and Senjanovic (1981); Magg and Wetterich (1980); Wetterich (1981); Schechter and Valle (1980); Foot et al. (1989); Ma (1998); Ma and Roy (2002),
@@
where $M$ is the mass of the heavy seesaw state and $v_{H}=246$ GeV is the Higgs vacuum expectation value (vev). A large $M$ naturally explains the smallness of active neutrino masses. A $U(1)_{B-L}$ symmetry provides a natural explanation for why $M$ is so large: the heavy seesaw masses are tied to the scale at which BLB-L is spontaneously broken, $v_{B-L}$, so if the scale of symmetry breaking is large, then so is the mass of the heavy seesaw state. Moreover, $U(1)_{B-L}$ is not always ad hoc, but appears in many Grand Unified Theory symmetry-breaking chains, giving further motivation for embedding the seesaw mechanism within this framework Gell-Mann et al. (1979); Buchmuller et al. (2005); Dror et al. (2020); Dunsky et al. (2022). The particle content for each type of seesaw is shown in Table 1.
@@ II.1 The Condition for Early Matter Domination
In this section, we derive the analytic conditions for right-handed neutrino early matter domination. We first present an approximate derivation, following the standard approach in the literature. While convenient, this estimate is inaccurate by roughly half an order of magnitude. We therefore provide a more accurate derivation, which can still be carried out analytically.
@@ II.1.1 Derivation I: Canonical derivation
As a concrete example, the Type-I seesaw introduces Standard Model singlets $N_{i}$ with degrees of freedom $g=\tfrac{7}{8}\times 2$. This gives Giudice et al. (1999)
@@
This illustrates how the onset of early matter domination in seesaw scenarios is directly controlled by the heavy state mass.
@@ II.1.2 Derivation II: Comoving Abundance Method
Once $T\ll M$, the species behaves as non-relativistic matter and its energy density is well approximated by $\rho_{N}=M\,Y\,s(T)$. In this regime, the average particle energy, averaging over momentum space, is $\langle E\rangle\simeq M+\mathcal{O}(T)$, so thermal corrections are negligible compared to the mass. Equating the right-handed neutrino and radiation energy densities $\rho_{N}=\rho_{R}$, sets the temperature for early matter domination
@@
where the superscript $i$ denotes the initial value. In what follows, we adopt Derivation II. For successful early matter domination by such particles, the following three conditions must be satisfied:
@@
    The particle must not have decayed by the time of early matter domination:
@@
We now investigate these three conditions for the $U(1)_{B-L}$ extended seesaw frameworks. If instead we consider non-thermal production channels, such as inflaton decay Asaka et al. (1999); Hahn-Woernle and Plümacher (2009); Ghoshal et al. (2024), curvaton decay Fong et al. (2023), generic and modulated sneutrino decay Mazumdar (2004); Allahverdi and Mazumdar (2003); Mazumdar and Perez-Lorenzana (2004); Afzal et al. (2025), Q-ball decay Fujii et al. (2001), from phase transition bubbles Dasgupta et al. (2022); Cataldi and Shakya (2024), preheating Cui et al. (2024) and reheating Ghoshal et al. (2023b), or primordial black holes evaporations Datta et al. (2021); Das et al. (2021); Perez-Gonzalez and Turner (2021); Barman et al. (2022); Bernal et al. (2022); Ghoshal et al. (2023a), only conditions (ii) and (iii) need to be imposed. To remain general, we shall leave the initial abundance $Y_{i}$ as a free parameter throughout our analysis.
@@ II.2.1 Type I: Global $U(1)_{B-L}$ case
The type I seesaw mechanism Minkowski (1977); Gell-Mann et al. (1979); Yanagida (1979) is among the most economical and widely studied explanations for the origin of light neutrino masses. In this framework, heavy right-handed neutrinos $N_{i}$ are introduced, which couple to the Standard Model lepton doublets $L_{\alpha}$ and the Higgs doublet $H$ through Yukawa interactions. After electroweak symmetry breaking, these interactions generate a Dirac mass term for the neutrinos. If the right-handed neutrinos also acquire large Majorana masses, the light neutrinos obtain naturally small masses via the seesaw mechanism. Since the Majorana mass term explicitly violates BLB-L by two units, the global $U(1)_{B-L}$ symmetry requires the introduction of an additional complex scalar field $\Phi$, carrying non-zero BLB-L charge Fu et al. (2023). When $\Phi$ acquires a vacuum expectation value, the BLB-L symmetry is broken, the right-handed neutrinos obtain Majorana masses, and the seesaw mechanism is realized. The Lagrangian is augmented by Yukawa, Majoron and Higgs portal interactions, as well as a quartic potential for the symmetry-breaking scalar
@@
$\Gamma_{\text{prod}}\sim\left(\frac{\lambda_{\phi}\ v_{B-L}\ \tilde{y}_{i}}{m_{\phi}^{2}}\right)^{2}T,\quad\Rightarrow\quad\frac{\Gamma_{\rm prod}}{H}\sim\left(\frac{\lambda_{\phi}\ v_{B-L}\ \tilde{y}_{i}}{m_{\phi}^{2}}\right)^{2}\ \frac{M_{\text{Pl}}}{T}\ .$ (13)
@@
$\frac{\Gamma_{\rm prod}}{H}\propto\frac{1}{T^{5}}$ (14)